x * (dy)/(dx) = x ^ 2 + 5y

1
Expert's answer
2021-12-21T12:16:31-0500

"y'-5y\/x=x"


for

"y'-5y\/x=0" :

"dy\/y=5dx\/x"

"lny_1=5lnx+lnc_1"

"y_1=c_1x^5"


then:

"y=x^5c(x)"

"5x^4c(x)+x^5c'(x)-5x^4c(x)=x"

"c'(x)=1\/x^4"

"c(x)=C-\frac{1}{3x^3}"


"y=(C-\frac{1}{3x^3})x^5"